# Sharp Error-Rate Transitions in Quantum QC-LDPC Codes under Joint BP Decoding

### v1

@@ Abstract
In this study, we report that quantum quasi-cyclic low-density parity-check codes decoded via joint belief propagation exhibit steep error-rate curves, despite the presence of error floors. To the best of our knowledge, this is the first observation of such threshold-like behavior for quantum codes with non-vanishing coding rate, excluding those decoded with non-binary BP decoders. Moreover, we find that dominant error events contributing to the error floor typically involve only a small number of bits. These findings suggest that the error floor is caused by trapping sets—specific subgraph structures in the Tanner graph—and indicate that identifying and avoiding such structures may lead to further reduction of the error floor.
@@ I   Introduction
A quantum error correction scheme exhibiting a steep frame error rate curve approaching the hashing bound and a low error floor has been reported in [1]. This decoding method utilizes a non-binary low-density parity-check (LDPC) decoder to decode quantum LDPC codes with computational complexity proportional to the code length. However, the decoding complexity scales quadratically with the size of the finite field. To the best of the author’s knowledge, except for the code proposed in [1], no quantum error-correcting codes with non-vanishing coding rates have been reported to exhibit steep performance curves. In contrast, threshold phenomena have been investigated for codes with vanishing rates [2]. This highlights a fundamental difference in decoding behavior between finite-rate and vanishing-rate quantum codes, suggesting that further analysis is needed to understand the mechanism behind steep transitions in the former.
One notable exception is [3], which demonstrates that a particular class of LDPC codes achieves a steep error-performance curve over a quantum erasure channel. While the channel model may not be of immediate practical relevance, the result provides a valuable theoretical reference and benchmark for future studies.

### v2

@@ Abstract
In this study, we report that quantum quasi-cyclic low-density parity-check codes decoded via joint belief propagation (BP) exhibit steep error-rate curves, despite the presence of error floors. To the best of our knowledge, this is the first observation of such threshold-like behavior for quantum LDPC codes with non-vanishing coding rate, excluding those decoded with non-binary BP decoders. Moreover, we find that dominant error events contributing to the error floor typically involve only a small number of bits. These findings suggest that the error floor is caused by trapping sets—specific subgraph structures in the Tanner graph—and indicate that identifying and avoiding such structures may lead to further reduction of the error floor.
@@ I   Introduction
A quantum error correction scheme exhibiting a steep frame error rate curve approaching the hashing bound and a low error floor has been reported in [1]. This decoding method utilizes a non-binary low-density parity-check (LDPC) decoder to decode quantum LDPC codes with computational complexity proportional to the code length. However, the decoding complexity scales quadratically with the size of the finite field.
Classical coding theory is well established, and efficient coding schemes have been developed that achieve theoretical limits. These schemes exhibit steep reductions in error rate and low error floors as the code length increases.
To the best of the authors’ knowledge, apart from the code proposed in [1], no quantum LDPC codes with non-vanishing coding rates have been reported to exhibit steep scaling behavior. In contrast, threshold phenomena have been extensively studied for quantum codes with vanishing rates [2], underscoring a fundamental distinction in decoding behavior between finite-rate and vanishing-rate quantum codes. This disparity suggests the need for further analysis to elucidate the mechanisms behind sharp transitions in the finite-rate regime. One related exception is [3], which demonstrates that a certain class of LDPC codes achieves steep performance curves over the quantum erasure channel. While the channel model may not directly reflect realistic noise, the result provides a valuable theoretical benchmark. Another notable example is [4], which reports that a rate-1/9 quantum turbo code exhibits scalable waterfall behavior with sharp error-rate transitions, further supporting the feasibility of such phenomena beyond LDPC-based constructions.